# Magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances in hot and dense strange hadronic matter

### v1

@@ II Chiral Quark Mean Field Model
The above expression is introduced in chiral effective models to account for the non-vanishing pseudoscalar meson masses and it satisfies the partial conserved axial-vector current relations for $\pi$ and $K$ mesons fortyfour ; fortysix ; fortyseven . The confinement of constituent quarks inside the baryons, due to the presence of a confining scalar-vector potential $\chi_{c}$, is introduced through the last term in the Lagrangian density, in eq. 1, and is written as
${\cal L}_{c}=-\bar{\psi}\chi_{c}\psi,$ (19)
where the confining potential $\chi_{c}$ is given by
@@
In eq. 20 above the coupling constant $k_{c}$ is taken to be $98\,\text{MeV}.\text{fm}^{-2}$.
In order to investigate the properties of asymmetric strange matter at finite temperature and density, we are using mean field approximation fortyfour . Under the influence of meson mean field, we express the Dirac equation for the quark field $\Psi_{qi}$ as
$\left[-i\vec{\alpha}\cdot\vec{\nabla}+\chi_{c}(r)+\beta m_{q}^{*}\right]\Psi_{qi}=e_{q}^{*}\Psi_{qi}.$ (21)
Here the quark $q$ ($q=u,d,s$) is in a baryon of type $i$ ($i=N,\Lambda,\Sigma,\Xi$)  and $\vec{\alpha}$ , $\beta$  are the usual Dirac matrices. The effective quark mass $m_{q}^{*}$ is defined as
$m_{q}^{*}=-g_{\sigma}^{q}\sigma-g_{\zeta}^{q}\zeta-g_{\delta}^{q}I^{3q}\delta+m_{q0},$ (22)
At zero baryon density, the masses of constituent quarks $u$, $d$, $s$ are expressed in terms of vacuum expectation values $\sigma_{0}$ and $\zeta_{0}$ of meson scalar fields $\sigma$ and $\zeta$, respectively. For the light $u$ and $d$ quarks, we have
$m_{u}=m_{d}=-g_{\sigma}^{q}\sigma_{0}=-\frac{g_{s}}{\sqrt{2}}\sigma_{0}.$ (23)
The values of coupling constant $g_{s}$ and confining parameter $k_{c}$ are fitted to obtain binding energy of $-16.0$ MeV at nuclear saturation density, $\rho_{0}=0.16$ $\text{fm}^{-3}$. This fit provides the value of effective mass of light quarks to be equal to $256$ MeV. Here $m_{q0}$ is zero for non-strange $u$ and $d$ quarks, whereas for strange $s$ quark $m_{q0}=\Delta m=77$ MeV. The term $\Delta m$ is introduced through the Lagrangian density ${\cal L}_{\Delta m}=-\Delta m\bar{\psi}m_{1}\psi$, where $S_{1}=\frac{1}{3}\left(1-\lambda_{8}\sqrt{3}\right)$. In eq. 21, $e_{q}^{*}$ in the Dirac equation is the effective energy of quarks and is expressed as fortyfour
@@
The in-medium mass of a baryon is expressed in terms of its effective energy $E^{*}_{i}$ and spurious center of mass momentum $p_{icm}$ as in Barik1985 ; Barik2013 ,
$M^{*}_{i}=\sqrt{E^{*2}_{i}-<p^{*2}_{icm}>}.$ (25)
Further the in-medium energies of $i^{th}$ baryon can be related to the in-medium energies of quarks and number of q type constituent quarks $n_{qi}$ in $i^{th}$ baryon by the expression
@@
Here, $E_{ispin}$ is a correction term in the effective energy of a baryon that accounts for spin-spin interaction between constituent quarks. The $E_{ispin}$ is determined so as to obtain the correct baryon masses at zero density. Finally, the thermodynamic potential $\Omega$ describing the isospin asymmetric strange hadronic matter at finite temperature and baryon density is written as fortyfour
$\Omega=-\frac{k_{B}T}{(2\pi)^{3}}\sum_{i}\gamma_{i}\int_{0}^{\infty}d^{3}k\left\{\ln\left(1+e^{-\left[E_{i}^{*}(k)-\nu_{i}^{*}\right]/k_{B}T}\right)+\ln\left(1+e^{-\left[E_{i}^{*}(k)+\nu_{i}^{*}\right]/k_{B}T}\right)\right\}-\mathcal{L}_{M}-\mathcal{V}_{\text{vac }},$ (27)
where the mesonic Lagrangian is $\mathcal{L}_{M}=\mathcal{L}_{VV}+\mathcal{L}_{\Sigma\Sigma}+\mathcal{L}_{\chi SB}$ and we have $E^{*}_{i}(k)=\sqrt{M_{i}^{*2}+k^{2}}$. In the above equation, the degeneracy factor is $\gamma_{i}=2$ for baryons and the summation is over hyperons and nucleons. In eq. 27, the effective chemical potential $\nu_{i}^{*}$ is expressed in terms of the usual baryon chemical potential $\mu_{i}$ as fortyfour ; fiftythree ; fiftyfour ; sixtytwo
@@
From eq. 27 the equations of motion of scalar and vector meson fields are calculated by minimizing the effective thermodynamic potential as (see APPENDIX)
@@
Equation 29 is solved for finite values of temperature and density of the medium. In order to account for isospin asymmetry in the hadronic medium a parameter $i_{a}=-\frac{\Sigma_{i}I_{3i}\rho_{i}}{\rho_{B}}$ is introduced, where $I_{3i}$ represents the third component of isopin and $\rho_{B}$ is baryonic density of the medium fiftyfour . To take into account hyperons in the medium, we define strangeness fraction $f_{s}=\frac{\Sigma_{i}\left|s_{i}\right|\rho_{i}}{\rho_{B}}$, here $|s_{i}|$ denotes the number of $s$ quarks in $i^{th}$ baryon.
@@
We have eq. 25 which addresses our aim to compute the effective baryon masses, by taking into account the effective constituent quark masses, baryon density and the spurious center of mass motion. In the upcoming section we shift the trajectory of our discussion towards the calculation of effective baryonic magnetic moments, by incorporating the medium modified constituent quark masses and the effective baryon masses.
III Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
@@
In the present section, methodology used to calculate the magnetic moments of ${\frac{1}{2}}^{-}$ baryon resonances has been discussed. The formalism of chiral constituent quark model ($\chi\mathrm{CQM}$) thirtyone , has been used to calculate the magnetic moment of the baryons through the following fluctuation process eight ; fiftyfive
@@
$q^{\uparrow\downarrow}\rightarrow\mathrm{GB}+q^{\prime\downarrow\uparrow}\rightarrow{\left(q\overline{q}^{\prime}\right)+q^{\prime\downarrow\uparrow}},$ (30)
@@
where GB represents the emitted Goldstone boson and $q\overline{q}^{\prime}+q^{\prime}$ constitute the sea quarks nineteen ; twenty ; twentyone ; twentytwo ; twentythree . The $\chi\mathrm{CQM}$ model incorporates the concept of confinement and chiral symmetry breaking twelve ; fiftysix , where the massless quarks acquire mass through spontaneous breaking of chiral symmetry twelve . In this approach we consider the emission of a GB, which further splits into a $q\overline{q}^{\prime}$ pair. The coupling between the quark and GBs is described by the effective interaction Lagrangian,
@@
$\mathcal{L}=g_{8}\overline{\mathbf{q}}\left(\Phi+\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}I\right)\mathbf{q}=g_{8}\overline{\mathbf{q}}\left(\Phi^{\prime}\right)\mathbf{q}.$ (31)
@@
This quark-GB coupling represented in this Lagrangian is weak enough to consider the fluctuation in eq. 30 as a small perturbation and the higher order fluctuations can therefore be neglected. The process of emission of GB from the quark gives rise to the probability of a spin-flip of the quark. In eq. 31, $\zeta^{\prime}=g_{1}/g_{8},$ where $g_{1}$ and $g_{8}$ denote the coupling constants for the singlet and octet GBs, respectively and $I$ is a $3\times 3$ identity matrix. The matrix of the GBs can be written as
@@
The $SU(3)$ chiral symmetry breaking is introduced in $\chi\mathrm{CQM}$ by considering that $m_{s}>m_{u,d}$ and the masses of GBs as non-degenerate $\left(m_{\eta^{\prime}}>m_{K,\eta}>m_{\pi}\right)$ fiftyseven . The parameter $a=\left|g_{8}\right|^{2}$ is denoted as the transition probability of chiral fluctuation $u(d)\rightarrow$ $d(u)+\pi^{+(-)}$, whereas $a\alpha^{2},a\beta^{2}$ and $a\zeta^{\prime 2}$ denote the probabilities of transitions of $u(d)\rightarrow s+K^{-(0)},u(d,s)\rightarrow$ $u(d,s)+\eta$ and $u(d,s)\rightarrow u(d,s)+\eta^{\prime}$, respectively nineteen ; twenty ; twentyone ; twentytwo ; twentythree . By thoroughly analysing the spin and flavor distribution functions of proton, the best fitted set of parameters crucial to our calculations are nineteen ; twentythree ; twentyfive ; twentyeight ; twentynine ; thirty ; thirtyone : $a=0.12,\alpha=\beta=0.45,\eta=-0.15$.
@@
Supposedly B ¿ is a SU(6) spin-flavor wave function of a baryon, then we have
$\left\langle B\left|\mu^{*S}_{B}\right|B\right\rangle=\Delta u\mu^{*}_{u}+\Delta d\mu^{*}_{d}+\Delta s\mu^{*}_{s},$ (34)
$\left\langle B\left|\mu^{*L}_{B}\right|B\right\rangle=\Delta u^{(1)}\mu^{*}_{u}+\Delta d^{(1)}\mu^{*}_{d}+\Delta s^{(1)}\mu^{*}_{s}.$ (35)
Here, $\mu^{*}_{q}(q=u,d,s)$ is the effective quark magnetic moment. The spin and orbital angular momentum polarizations are quantitatively comprehended from $\Delta q=q^{\uparrow}-q^{\downarrow}$ and $\Delta q^{(1)}=q^{(1)}-q^{(-1)}$, respectively. The polarizations can be calculated using the number operator $\mathcal{N}$ (=$\mathcal{N}^{\uparrow\downarrow}+\mathcal{N}^{\pm 1}$) as
$\widehat{B}\equiv\langle B|\mathcal{N}|B\rangle.$ (36)
Here $\mathcal{N}^{\uparrow\downarrow}$ and $\mathcal{N}^{(\pm 1)}$ of the number operator $\mathcal{N}$ are expressed as
$\mathcal{N}^{\uparrow\downarrow}=$ $n_{u\uparrow}u^{\uparrow}+n_{u\downarrow}u^{\downarrow}+n_{d\uparrow}d^{\uparrow}+n_{d^{\downarrow}}d^{\downarrow}+n_{s^{\uparrow}}s^{\uparrow}+n_{s\downarrow}s^{\downarrow},$ (37)
$\mathcal{N}^{(\pm 1)}=$ $n_{u^{(1)}}u^{(1)}+n_{u(-1)}u^{(-1)}+n_{d^{(1)}}d^{(1)}+n_{d^{(-1)}}d^{(-1)}+n_{s^{(1)}}s^{(1)}+n_{s^{(-1)}}s^{(-1)},$ (38)
with $n_{q^{\uparrow}}\left(n_{q^{\downarrow}}\right)$ denoting the number of spin up (down) quarks and $n_{q^{(1)}}\left(n_{q^{(-1)}}\right)$ being the number of quarks with the orbital angular momentum projection $m_{L}=1$ $\left(m_{L}=-1\right)$.
Following the Cheng Li mechanism ten , the calculation of total baryonic magnetic moment involves the contribution of valence and sea quark spin polarizations and orbital angular momentum of sea quarks. The spin part in eq. 33, can be obtained from contributions due to valence, sea and orbital angular momentum of sea quarks and is given by
@@
In eq. 40, $P_{[q,GB]}$ is the probability of emission of GBs from a quark $q^{\uparrow(\downarrow)}$ fiftyeight given by
$P(u\rightarrow\mathrm{GB})=P(d\rightarrow\mathrm{GB})=\frac{a}{6}\left(9+6\alpha^{2}+\beta^{2}+2\zeta^{\prime 2}\right),$ (41)
$P(s\rightarrow\mathrm{GB})=\frac{a}{3}\left(6\alpha^{2}+2\beta^{2}+\zeta^{\prime 2}\right),$
@@
and $\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}$ represents the probability of transforming a $q^{\uparrow(\downarrow)}$ quark and it is given by nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; fiftythree ; fiftyfour
$\left|\psi\left(u^{\uparrow(\downarrow)}\right)\right|^{2}=\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)u^{\uparrow(\downarrow)}+ad^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$ (42)
$\left|\psi\left(d^{\uparrow(\downarrow)}\right)\right|^{2}=au^{\uparrow(\downarrow)}+\frac{a}{6}\left(3+\beta^{2}+2\zeta^{\prime 2}\right)d^{\uparrow(\downarrow)}+a\alpha^{2}s^{\uparrow(\downarrow)},$
$\left|\psi\left(s^{\uparrow(\downarrow)}\right)\right|^{2}=a\alpha^{2}u^{\uparrow(\downarrow)}+a\alpha^{2}d^{\uparrow(\downarrow)}+\frac{a}{3}\left(2\beta^{2}+\zeta^{\prime 2}\right)s^{\uparrow(\downarrow)}.$
The orbital angular momentum contribution due to sea quarks, is expressed as ten ; fourteen ; twentyeight ; twentynine ; thirty ; thirtyone
@@
$\mu^{*}\left(q^{\uparrow}\rightarrow q^{\prime\downarrow}\right)=\frac{e_{q}}{2m^{*}_{q}}\left\langle l_{q}\right\rangle+\frac{e_{q}-e_{q\prime}}{2M_{\mathrm{GB}}}\left\langle l_{\mathrm{GB}}\right\rangle,$ (43)
@@
where $l_{q}\right>=\frac{M_{\mathrm{GB}}}{M_{\mathrm{GB}}+m^{*}_{q}}$ and $l_{\mathrm{GB}}\right>=\frac{m^{*}_{q}}{M_{\mathrm{GB}}+m^{*}_{q}}$. The quantities $\left(l_{q},l_{\mathrm{GB}}\right)$ and $\left(m^{*}_{q},M_{\mathrm{GB}}\right)$ stands for the orbital angular momenta and masses of quarks and GBs, respectively. The orbital moment of each process in eq. 43, when multiplied by the probability of occurrence of the process, produces the magnetic moment due to all possible transitions for a given valence quark. This can be mathematically expressed as,
${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(u_{\uparrow}\rightarrow u_{\downarrow}\right)\right.$ (44)
$\left.+\mu^{*}\left(u_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(u_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\mu^{*}\left(d_{\uparrow}\rightarrow u_{\downarrow}\right)+\left(\frac{1}{2}+\frac{\beta^{2}}{6}+\frac{\zeta^{\prime 2}}{3}\right)\right.$
×$\left.\times\mu^{*}\left(d_{\uparrow}\rightarrow d_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(d_{\uparrow}\rightarrow s_{\downarrow}\right)\right],$
${\left[\mu^{*}\left(s_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow u_{\downarrow}\right)+\alpha^{2}\mu^{*}\left(s_{\uparrow}\rightarrow d_{\downarrow}\right)\right.$
$\left.+\left(\frac{2}{3}\beta^{2}+\frac{\zeta^{\prime 2}}{3}\right)\mu^{*}\left(s_{\uparrow}\rightarrow s_{\downarrow}\right)\right].$
@@
The orbital moments of the $u,d$ and $s$ quarks can also be expressed in terms of the $\chi\mathrm{CQM}$ parameters $(a,\alpha,\beta,\zeta^{\prime})$, masses of GBs $\left(M_{\pi},M_{\eta\prime},M_{K},M_{\eta}\right)$ and effective quark masses $\left(m^{*}_{u},m^{*}_{d},m^{*}_{s}\right)$, respectively as thirtyone ; fiftythree ; fiftyfour :
${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\frac{3m_{u}^{*2}}{2M_{\pi}\left(m_{u}^{*}+M_{\pi}\right)}-\frac{\alpha^{2}\left(M_{K}^{2}-3m_{u}^{*2}\right)}{2M_{K}\left(m_{u}^{*}+M_{K}\right)}\right.+\frac{\beta^{2}M_{\eta}}{6\left(m_{u}^{*}+M_{\eta}\right)}$
$\left.+\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{3\left(m_{u}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{u},$
${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{3\left(M_{\pi}^{2}-2m_{d}^{*2}\right)}{4M_{\pi}\left(m^{*}_{d}+M_{\pi}\right)}-\frac{\alpha^{2}M_{K}}{2\left(m_{d}^{*}+M_{K}\right)}-\frac{\beta^{2}M_{\eta}}{12\left(m_{d}^{*}+M_{\eta}\right)}\right.$ (45)
$\left.-\frac{\zeta^{\prime 2}M_{\eta^{\prime}}}{6\left(m_{d}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{d},$

### v2

@@ 2.1 Chiral Quark Mean Field Model
The above expression accounts for the non-zero masses of pseudoscalar mesons and agrees well with the partial conserved axial-vector current relations of $\pi$ and $K$ mesons fortyfour ; fortysix ; fortyseven . Unlike the models which uses the concept of spherical-bag boundary, like MIT bag Model, the CQMF model treats the constituent quarks to be moving inside the baryons forming a bounded system, by virtue of a confining potential $\chi_{c}$. The use of an average confining potential with quarks moving around the baryon core, has been fruitful in improving the static baryonic properties like magnetic moment of proton Barik1985 . The last term in the Lagrangian density, in eq. 1, incorporates this as
${\cal L}_{cp}=-\bar{\psi}\chi_{c}\psi,$ (20)
and the confining potential $\chi_{c}$, is expressed as
@@
In eq. 21 above the coupling constant $k_{c}$ is considered to be $98\,\text{MeV}.\text{fm}^{-2}$.
In light of mean field approximation, we are to investigate the asymmetric strange matter properties at finite temperature and density, we are using fortyfour . In the presence of meson mean field, we express the Dirac equation for the quark field $\Psi_{qi}$ as
$\left[-i\vec{\alpha}\cdot\vec{\nabla}+\beta m_{q}^{*}+\chi_{c}(r)\right]\Psi_{qi}=e_{q}^{*}\Psi_{qi}.$ (22)
Here the quark $q$ is in a baryon of type $i$ and $\vec{\alpha}$ , $\beta$  are the Dirac matrices. By definition, the in-medium constituent quark mass $m_{q}^{*}$ is
$m_{q}^{*}=-g_{\sigma}^{q}\sigma-g_{\zeta}^{q}\zeta-g_{\delta}^{q}I^{3q}\delta+m_{q0}.$ (23)
In free space, the masses of constituent quarks $u$, $d$, $s$ are expressed in terms of $\sigma_{0}$ and $\zeta_{0}$, by the relation
$m_{u}=m_{d}=-\frac{g_{s}}{\sqrt{2}}\sigma_{0}=-g_{\sigma}^{q}\sigma_{0}.$ (24)
The coupling constant $g_{s}$ and confining parameter $k_{c}$ are fitted to obtain binding energy of $-16.0$ MeV at nuclear saturation density, $\rho_{0}=0.16$ $\text{fm}^{-3}$ fortyfive . This fit provides the value of effective mass of light quarks to be equal to $256$ MeV. Here $m_{q0}$ is zero for non-strange $u$ and $d$ quarks, while for $s$ quark $m_{q0}=\Delta m=77$ MeV. In eq. 22, $e_{q}^{*}$ in the Dirac equation is the effective energy of quarks and is expressed as fortyfour
@@
The in-medium mass of a baryon is expressed as,
$M^{*}_{i}=\sqrt{E^{*2}_{i}-<p^{*2}_{icm}>},$ (26)
in terms of its in-medium energy $E^{*}_{i}$ and spurious center mass motion $p_{icm}$ Barik1985 ; Barik2013 . Spuirous center of mass momentum $p_{icm}$ corrections are added in expression for $M_{i}^{*}$ to further improve the overlooked independent quark motion inside the baryons. The explicit expression for $<p^{*2}_{icm}>$ being summation over average value of square of each constituent quark momentum ($<p^{*2}_{cm}>_{q}$) Barik1985 , i.e.,
$<p^{*2}_{icm}>=\sum_{q}<p^{*2}_{cm}>_{q}=\sum_{q}\frac{11e_{q}^{*}+m_{q}^{*}}{6(3e_{q}^{*}+m_{q}^{*})}(e_{q}^{*2}-m_{q}^{*2}).$ (27)
As can be seen from the above equation, spurious center of mass motion corrections are density and temperature dependent. In Ref. fortyseven , spurious center of mass motion corrections are considered as independent of medium effects. Further the in-medium energies of $i^{th}$ baryon can be related to the in-medium energies of quarks and number of q type constituent quarks $n_{qi}$ in $i^{th}$ baryon by the expression
@@
Here, $E_{ispin}$ accounts for spin-spin interaction of constituent quarks and serves as a correction term in the effective energy of a baryon. The $E_{ispin}$ is computed so as to obtain the correct baryon masses at zero density. Finally, the thermodynamic potential $\Omega$ describing the isospin asymmetric strange hadronic matter at finite temperature and baryon density is written as fortyfour
$\Omega=-\frac{k_{B}T}{(2\pi)^{3}}\sum_{i}\gamma_{i}\int_{0}^{\infty}\left\{\ln\left(e^{-\left[E_{i}^{*}(k)-\nu_{i}^{*}\right]/k_{B}T}+1\right)+\ln\left(e^{-\left[E_{i}^{*}(k)+\nu_{i}^{*}\right]/k_{B}T}+1\right)\right\}d^{3}k-\mathcal{L^{\prime}}-\mathcal{V}_{\text{vac }},$ (29)
where the mesonic Lagrangian is $\mathcal{L^{\prime}}=\mathcal{L}_{VV}+\mathcal{L}_{\Sigma\Sigma}+\mathcal{L}_{\chi SB}$ and we have $E^{*}_{i}(k)=\sqrt{M_{i}^{*2}+k^{2}}$. In the above equation, the degeneracy factor is $\gamma_{i}=2$ for baryons and the summation is over hyperons and nucleons. In eq. 29, the effective chemical potential $\nu_{i}^{*}$ is expressed in terms of the usual baryon chemical potential $\mu_{i}$ as fortyfour ; fiftythree ; fiftyfour ; sixtytwo
@@
From eq. 29 the equations of motion of scalar and vector meson fields are calculated by minimizing the effective thermodynamic potential as
@@
Equation 31 is solved for finite values of temperature and densities of the medium. Note that the confining potential $\chi_{c}$ contribute to the effective energy $e_{q}^{*}$ and hence, masses $m_{q}^{*}$ of quarks through the solution of Dirac equation. The effective quark energies are used in calculations of baryon energy which further enter into the coupled system of equations for scalar and vector fields through the definition of scalar and vector density of baryons fortyfour . In order to account for isospin asymmetry in the hadronic medium a parameter $i_{a}=-\frac{\Sigma_{i}I_{3i}\rho_{i}}{\rho_{B}}$ is introduced, where $I_{3i}$ represents the third component of isopin and $\rho_{B}$ is baryonic density of the medium fiftyfour . To take into account hyperons in the medium, we define strangeness fraction $f_{s}=\frac{\Sigma_{i}\left|s_{i}\right|\rho_{i}}{\rho_{B}}$, here $|s_{i}|$ denotes the number of $s$ quarks in $i^{th}$ baryon. Note that in the present manuscript, the difference in the $\frac{1}{2}^{-}$ baryon resonances and positive parity octet baryons comes through the fitting of $E_{ispin}$ value to their respective vacuum masses. The calculations of effective masses of $\frac{1}{2}^{-}$ baryon resonances proceed in two steps: first we solve the system of equations corresponding to the scalar ($\sigma,\zeta$ and $\delta$) and vector fields ($\omega,\rho$ and $\phi$) in the strange hadronic medium (considering only positive parity octet baryons and utilizing Eqs. (26) and (28) for in-medium effective masses and energies) at finite density and temperature (see Eq. (31)). This gives us the density and temperature dependent values of scalar and vector fields which are used in the second step to obtain the effective masses of $\frac{1}{2}^{-}$ baryon resonances. For this, first we use the vacuum ($\rho_{B}=0,T=0,i_{a}=0,f_{s}=0$) values of fields calculated in first step, in Eqs. (26) and (28 and fit $E_{ispin}$ to vacuum masses of $\frac{1}{2}^{-}$ baryon resonances. Using these $E_{ispin}$ values and density and temperature dependent values of scalar and vector fields calculated in the first step (strange matter), the in-medium masses of $\frac{1}{2}^{-}$ baryon resonances are calculated using Eqs. (26) in the strange hadronic medium.
@@
We have eq. 26 which addresses our aim to compute the effective baryon masses, by taking into account the effective constituent quark masses, baryon density and the spurious center of mass motion. In the upcoming section we shift the trajectory of our discussion towards the calculation of effective baryonic magnetic moments, by incorporating the effective in-medium masses of constituent quarks and the baryons under study.
2.2 Magnetic Moment of ${\frac{1}{2}}^{-}$ baryon resonances
@@
The formalism of chiral constituent quark model ($\chi\mathrm{CQM}$) thirtyone , has been implemented to calculate the magnetic moment of the baryons through the following fluctuation process eight ; fiftyfive
@@
$q^{\uparrow\downarrow}\rightarrow\mathrm{NGB}+q^{\prime\downarrow\uparrow}\rightarrow{\left(q\overline{q}^{\prime}\right)+q^{\prime\downarrow\uparrow}},$ (32)
@@
where the emitted Nambu-Goldstone boson (NGB) furthermore splits into a $q\overline{q}^{\prime}$ pair, and $q\overline{q}^{\prime}+q^{\prime}$ constitute the sea quarks nineteen ; twenty ; twentyone ; twentytwo ; twentythree . In $\chi\mathrm{CQM}$ model the massless confined quarks obtain its mass by virtue of spontaneous chiral symmetry breaking twelve ; fiftysix . The coupling between the quark and NGBs is described by the effective interaction Lagrangian
@@
$\mathcal{L}=g_{8}\overline{\mathbf{q}}\left(\Phi^{\prime}\right)\mathbf{q}=g_{8}\overline{\mathbf{q}}\left(\zeta^{\prime}\frac{\eta^{\prime}}{\sqrt{3}}I+\Phi\right)\mathbf{q},$ (33)
@@
where $\zeta^{\prime}=g_{1}/g_{8},$ where $g_{1}$ and $g_{8}$ denote the coupling constants for the singlet and octet NGBs respectively and $I$ is an order 3 identity matrix. This quark-NGB coupling represented in this Lagrangian is quite feeble and hence, the fluctuation in eq. 32 is practically taken to be a minor perturbation. The spin-flip process is an outcome of the emission of NGB.
The NGB matrix is defined by
@@
The $SU(3)$ chiral symmetry breaking is introduced in $\chi\mathrm{CQM}$ by treating the masses of NGBs as non-degenerate and considering that $m_{s}>m_{u,d}$ fiftyseven . The parameter for the transition probability of chiral fluctuations: $u(d)\rightarrow$ $d(u)+\pi^{+(-)}$ is denoted by $a(=\left|g_{8}\right|^{2})$, $u(d)\rightarrow s+K^{-(0)}$ by $a\alpha^{2}$, $u(d,s)\rightarrow$ $u(d,s)+\eta$ by $a\beta^{2}$ and $u(d,s)\rightarrow u(d,s)+\eta^{\prime}$ is represented by $a\zeta^{\prime 2}$ nineteen ; twenty ; twentyone ; twentytwo ; twentythree . By thoroughly analyzing the spin and flavor distribution functions of proton, the best fitted set of parameters crucial to our calculations are nineteen ; twentythree ; twentyfive ; twentyeight ; twentynine ; thirty ; thirtyone : $a=0.12,\alpha=\beta=0.45,\zeta^{\prime}=-0.15$.
@@
Following the Cheng Li mechanism ten , the calculation of total baryonic magnetic moment involves the valence and sea quark contributions and the orbital angular momentum of sea quarks. The spin part in eq. 35, comprises of these individual contributions and is given by
@@
In the above expression, the emission probability of NGBs from a quark $q^{\uparrow(\downarrow)}$ is $P_{[q,NGB]}$ which can be fetched from ref. fiftyeight , and $\left|\psi\left(q^{\uparrow(\downarrow)}\right)\right|^{2}$ represents the probability of transforming a $q^{\uparrow(\downarrow)}$ quark nineteen ; twenty ; twentyone ; twentytwo ; twentythree ; fiftythree ; fiftyfour . The orbital angular momentum contribution due to sea quarks, is given by ten ; fourteen ; twentyeight ; twentynine ; thirty ; thirtyone
@@
$\mu^{*}\left(q^{\uparrow}\rightarrow q^{\prime\downarrow}\right)=\frac{e_{q}}{2m^{*}_{q}}\left\langle l_{q}\right\rangle+\frac{e_{q}-e_{q\prime}}{2M_{\mathrm{NGB}}}\left\langle l_{\mathrm{NGB}}\right\rangle,$ (38)
@@
where $l_{q}\right>=\frac{M_{\mathrm{NGB}}}{M_{\mathrm{NGB}}+m^{*}_{q}}$ and $l_{\mathrm{NGB}}\right>=\frac{m^{*}_{q}}{M_{\mathrm{NGB}}+m^{*}_{q}}$. The quantities $\left(l_{q},m^{*}_{q}\right)$ and $\left(l_{\mathrm{NGB}},M_{\mathrm{NGB}}\right)$ stands for the mass and orbital angular moment of constituent quarks and NGBs, respectively. The probability of each process to occur, allows us to compute the magnetic moment due to all possible transitions for a given valence quark.
@@
The orbital moments of the constituent quarks can also be written using the $\chi\mathrm{CQM}$ parameters $(a,\alpha,\beta,\zeta^{\prime})$, NGB masses $\left(M_{\pi},M_{\eta\prime},M_{K},M_{\eta}\right)$ and effective quark masses $\left(m^{*}_{u},m^{*}_{d},m^{*}_{s}\right)$, respectively as thirtyone ; fiftythree ; fiftyfour :
${\left[\mu^{*}\left(u_{\uparrow}\rightarrow\right)\right]=}$ $a\left[\frac{3}{2}\frac{m_{u}^{*2}}{M_{\pi}\left(m_{u}^{*}+M_{\pi}\right)}-\frac{\alpha^{2}}{2}\frac{\left(M_{K}^{2}-3m_{u}^{*2}\right)}{M_{K}\left(M_{K}+m_{u}^{*}\right)}\right.+\frac{\beta^{2}}{6}\frac{M_{\eta}}{6\left(M_{\eta}+m_{u}^{*}\right)}$
$\left.+\frac{\zeta^{\prime 2}}{3}\frac{M_{\eta^{\prime}}}{\left(m_{u}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{u},$
${\left[\mu^{*}\left(d_{\uparrow}\rightarrow\right)\right]=}$ $-2a\left[\frac{3}{4}\frac{\left(M_{\pi}^{2}-2m_{d}^{*2}\right)}{M_{\pi}\left(m^{*}_{d}+M_{\pi}\right)}-\frac{\alpha^{2}}{2}\frac{M_{K}}{\left(m_{d}^{*}+M_{K}\right)}-\frac{\beta^{2}}{12}\frac{M_{\eta}}{\left(m_{d}^{*}+M_{\eta}\right)}\right.$ (39)
$\left.-\frac{\zeta^{\prime 2}}{6}\frac{M_{\eta^{\prime}}}{\left(m_{d}^{*}+M_{\eta^{\prime}}\right)}\right]\mu^{*}_{d},$